OpenStudy (goformit100):

How to do it mentally John has 800 dollars in his checking account. He received from his job a check for 1,200 dollars and deposit the amount in his checking account. How much money does he have in his checking after the deposit?

OpenStudy (anonymous):

800 + 1200 = 2000 (0 + 0 = 0, 0+0 = 0, 2 + 8 = 10, carry the one over, 1 + 0 + 1 = 200) 2000.

OpenStudy (goformit100):

Thankyou Madm